OpenStudy (anonymous):

What is the solution of the system of equations? -3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9

OpenStudy (anonymous):

-3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9 multiply the first eq by -1 3x+4y+3z=7 2x-6y+2z=3 5x-2y+5z=9 multiply the first with 2 and the second with 3 6x+8y+6z=14 6x-18y+6z=9 5x-2y+ 5z = 9 subtract second eq from the first 6x+8y+6z=14 26y=5 => y = 5/26 5x-2y+ 5z = 9 substitute y in eq 1 and 3 3x+20/26+3z=14 y = 5/26 5x- 10/26+ 5z = 9 multiply the first by 5 and the third eq by 3 15x+100/26+15z=70 y = 5/26 15x- 30/26+ 15z = 27 substract the third eq from the first 130/26 = 43 which is false so the system has no solution